OpenStudy (anonymous):

Which rectangular equation is equivalent to the polar equation r cos theta=-2? A. x= -2x^2 B. y=-2 C. x=2 D.x^2+ y^2=-2?

OpenStudy (anonymous):

\[r \cos \theta = -2\] note that \[x = rcos \theta\]\[y= r \sin \theta\]\[r^2=x^2+y^2\]

OpenStudy (anonymous):

c. x=-2?************

OpenStudy (anonymous):

Yes :)

OpenStudy (triciaal):

was that a typo true x = -2 but C says x = 2

OpenStudy (anonymous):

Yup

OpenStudy (triciaal):